# DPBridge: Latent Diffusion Bridge for Dense Prediction

### v3

@@ Appendix C Network Architecture
The basic structure of our network follows the original design of Stable Diffusion. To better utilize visual information ($\mathbf{z}_{T}$ in our case), we concatenating both intermediate samples and input image along the feature dimension and send in the noise prediction network as a whole. To address the dimension incompatibility with pretrained backbone, the input channels of the network are thus doubled to accommodate the concatenated input. In addition, as suggested by [28], we duplicate the weight tensor of these accommodated layers and divide the values by two to preserve the input magnitude to subsequent layers. As DPBridge is an image-conditioned generative framework, we employ null-text embedding for all cross-attention layers to eliminate the impact of text prompts.
@@ Appendix D Implementation Details
During training, the number of timesteps of is set to be 1000, and we use 50 sampling steps during inference with considerations of both quality and efficiency. Training our method takes 30K iterations using the Adam optimizer with a base learning rate of $3\times 10^{-5}$ and an exponential decay strategy after 100-step warm-up. The batch size is 2, but with 8 steps gradient accumulation, the effective batch size is equivalent to 16. We also apply random horizontal flipping augmentation during training. The whole process takes approximately 3 days to complete on a single Nvidia RTX 3090Ti GPU.
@@ E.1 Additional Quantitative Results
In this section, we further evaluate the robustness of DPBridge system by perturbing the input image with different types and different level intensities of noise. In this experiment setting, we use Stable Diffusion 1.5 as pretrained backbone. The evaluation is conducted using under both depth estimation and surface normal prediction tasks. We use NYUv2 and KITTI to report the depth estimation results, as for surface normal prediction, we report the results on NYUv2 and iBims-1. Table 4 reveals that the model’s performance degrades as noise intensity increases, but the model performance remains acceptable under moderate-level-intensity noise. Among noise types, uniform noise shows the mildest effect on degradation, while salt & pepper and high-intensity gaussian noise cause the most significant performance decline. In addition, the noise intensity has more impact on surface normal prediction than depth estimation, showcasing the inherent characteristics between different tasks. These results emphasize the model’s robustness under moderate noise but still expose its vulnerability to extreme or specific types of noise.
@@ E.2 Additional Qualitative Results
To further evaluate the effectiveness and generalization capability of DPBridge, we extend the framework to additional dense prediction tasks, including semantic segmentation, optical flow estimation, and style transfer. Qualitative results are shown in Figure 7, Figure 8, Figure 9, and Figure 10. These extensions are implemented by framing each task as a continuous-valued image-to-image translation problem, and replacing the paired ground truth data accordingly. For semantic segmentation, we convert discrete labels into color maps and train the bridge model to predict colors in RGB space rather than class indices. For optical flow, we define the bridge process between the flow image and the first frame of the input RGB pair. To accommodate this, we modify the UNet architecture to condition on both RGB frames during inference and represent flow as a color-encoded image, enabling prediction in RGB space rather than raw uv coordinates. Across all tasks, DPBridge produces visually compelling results, demonstrating its versatility as a unified framework that can be easily adapted to a range of dense prediction scenarios. We note that these results are qualitative; quantitative evaluation and task-specific architectural refinements are left for future work.
Figure 7: Qualitative examples of applying DPBridge to semantic segmentation task on CityScape [12] dataset. We transform discrete labels into color maps and us our bridge model to predict colors in RGB space instead of categorical labels.
Figure 8: Qualitative examples of applying DPBridge to semantic segmentation task on CelebAMask-HQ [30] dataset. We transform discrete labels into color maps and us our bridge model to predict colors in RGB space instead of categorical labels.
Figure 9: Qualitative examples of applying DPBridge to optical flow prediction task on MPI-Sintel [5] dataset. The input image is the overlayed RGB pair, and we transform the flow map into RGB space so that the diffusion bridge process can be constructed accordingly.
Figure 10: Qualitative examples of applying DPBridge to style transfer task on Face-to-Comics [48] dataset.
Appendix F Broader Impacts
DPBridge aims to advance the field of dense prediction through a generative modeling perspective, integrating the strengths of diffusion bridge formulation with visual priors from pretrained backbones. Our proposed method is a unified framework that can be easily extend to different tasks, including depth prediction, surface normal estimation, semantic segmentation, optical flow, etc., which will benefit the whole dense prediction research community. In addition, improved generalization and precision in dense prediction can also benefit a wide range of real-world applications, such as autonomous driving, augmented reality, and digital content creation.

### v4

@@ Appendix C Network Architecture
The basic structure of our network follows the original design of Stable Diffusion. To better utilize visual information ($\mathbf{z}_{T}$ in our case), we concatenating both intermediate samples and input image along the feature dimension and send in the noise prediction network as a whole. To address the dimension incompatibility with pretrained backbone, the input channels of the network are thus doubled to accommodate the concatenated input. In addition, as suggested by [ke2024repurposing], we duplicate the weight tensor of these accommodated layers and divide the values by two to preserve the input magnitude to subsequent layers. As DPBridge is an image-conditioned generative framework, we employ null-text embedding for all cross-attention layers to eliminate the impact of text prompts.
@@ Appendix D Implementation Details
During training, the number of timesteps of is set to be 1000, and we use 50 sampling steps during inference with considerations of both quality and efficiency. Training our method takes 30K iterations using the Adam optimizer with a base learning rate of $3\times 10^{-5}$ and an exponential decay strategy after 100-steps warm-up. The batch size is 2, but with 8 steps gradient accumulation, the effective batch size is equivalent to 16. We also apply random horizontal flipping augmentation during training. The whole process takes approximately 3 days to complete on a single Nvidia RTX 3090Ti GPU.
@@ E.1 Additional Quantitative Results
In this section, we further evaluate the robustness of DPBridge system by perturbing the input image with different types and different levels of intensity of noise. In this experiment setting, we use Stable Diffusion 1.5 as pretrained backbone. The evaluation is conducted under both depth estimation and surface normal prediction tasks. We use NYUv2 and KITTI to report the depth estimation results, and for surface normal prediction, we report the results on NYUv2 and iBims-1. Tab.˜4 reveals that the model’s performance degrades as noise intensity increases, but the model performance remains acceptable under moderate-level-intensity noise. Among noise types, uniform noise shows the mildest effect on degradation, while salt & pepper and high-intensity Gaussian noise cause the most significant performance decline. In addition, the noise intensity has more impact on surface normal prediction than depth estimation, showcasing the inherent characteristics between different tasks. These results emphasize the model’s robustness under moderate noise but still expose its vulnerability to extreme or specific types of noise.
@@ E.2 Additional Qualitative Results
To further evaluate the effectiveness and generalization capability of DPBridge, we extend the framework to additional dense prediction tasks, including semantic segmentation, optical flow estimation, edge detection, and style transfer. Qualitative results are shown in Figure˜7, Figure˜9, Figure˜8, and Figure˜10. These extensions are implemented by framing each task as a continuous-valued image-to-image translation problem and replacing the paired ground truth data accordingly. For semantic segmentation, we convert discrete labels into color maps and train the bridge model to predict colors in RGB space rather than class indices. For optical flow, we define the bridge process between the flow image and the first frame of the input RGB pair. To accommodate this, we modify the UNet architecture to condition on both RGB frames during inference and represent flow as a color-encoded image, enabling prediction in RGB space rather than raw uv coordinates. Across all tasks, DPBridge produces visually compelling results, demonstrating its versatility as a unified framework that can be easily adapted to a range of dense prediction scenarios. We note that these results are qualitative; quantitative evaluation and task-specific architectural refinements are left for future work.
Figure 7: Qualitative examples of applying DPBridge to semantic segmentation task on CityScape [cordts2016cityscapes] dataset. We transform discrete labels into color maps and us our bridge model to predict colors in RGB space instead of categorical labels.
Figure 8: Qualitative examples of applying DPBridge to optical flow prediction task on MPI-Sintel [butler2012naturalistic] dataset. The input image is the overlayed RGB pair, and we transform the flow map into RGB space so that the diffusion bridge process can be constructed accordingly.
Figure 9: Qualitative examples of applying DPBridge to edge detection task on MultiCue [mely2016systematic] dataset.
Figure 10: Qualitative examples of applying DPBridge to style transfer task on Face-to-Comics [face2comics] dataset.